### v3

@@ 7.2 Unveiling the Definitions of Good Data
What signifies the most a good instruction datapoint remains an open question.
Unfortunately, there exists no unified criteria on discriminating “good" instructions from “bad" ones. Essentially, the definitions on the general data “quality" differ from task to task and domain to domain Evans and Murshudov (2013); Flach (2012); Albalak et al. (2024). Although existing quality measurement methods can be categorized in terms of quality, diversity, and importance under the present study, they all exhibit more or less ad-hoc properties in methodology. First, studies on instruction tuning are often targeted at improving the performance of LLMs on downstream task. No matter whether these tasks are of general-purpose (e.g., common NLP tasks on leaderboard Myrzakhan et al. (2024); Wolf et al. (2019)) or domain-specific applications, such task-orientated data selection itself is only a “proxy" for exploring the underlying “quality" measurement. Especially for coreset sampling methods that directly employ the evaluation set or testing set for distribution matching or importance estimation, instructions that resemble the most to the testing set or bring about performance gains are judged as “good" data. However, such “good" data cannot be easily transferred to another LLM of completely different architecture and parameters. Each time the entire pipeline has to be enforced for a novel task, making it difficult to accumulate universally-acknowledged high-quality data for archiving. Second, each method has an individual quality evaluation system and very few of them ever tried to justify their design and interpret the philosophy behind. It is difficult to validate whether certain component of the selection pipeline can be replaced or removed for better serving a new task-of-interest.
Accordingly, further academic explorations include: 1) to present a more unified, generally applicable definitions on “good" instruction datapoints in terms of fine-grained aspects, and 2) to improve interpretability and explanability of the selection pipeline beyond empirical design.
@@
The three aspects we used to categorize data assessment methods are actually overlapping with each other, where the “boundary" between two measuring dimension is often hard to explicitly defined. Under such circumstance, the definition of good data can be perceived as the weighted, biased mixture of quality, diversity, and importance. Existing methods are not flexible in dynamically adjusting the mixing weights to adapt to different downstream tasks. Instead, their priority order of the three dimensions is implicitly encoded into the selection of instructions. For instance,  Liu et al. (2023b) emphasizes quality and importance equally by first establishing the relative ranking of all samples in both quality and complexity. The subset is formed by consecutively selecting the top-ranked samples in sequence, with diversity intervened via ruling out heavily homogenized examples. Such hard-coded, greedy treatment to quality, diversity, and importance is not applicable to scenarios where the behavior of LLMs is expected to cater to varied preference.
@@ 7.3 Scaling Up Datasets
In the analysis of the disadvantages in exploiting the entire instruction dataset for alignment, putting aside the issue of long training time, we notice that the performance of fine-tuning the entire dataset might not be the optimal. There often exists a critical point of the best selection proportion, and such proportion varies from dataset to dataset. When more instruction datasets from diverse domains and tasks are incorporated, it becomes more difficult to nail down the best selection proportion for three main reasons. First, a large proportion of noise exists in the open datasets and few noisy samples can already cause tremendous negative impact on performance Song et al. (2022). During the pre-processing of instruction dataset, noise can be unintentionally introduced in instruction preparation (e.g., missing context or system prompt), response generation (e.g., unverified or mismatched answers), format wrapping (e.g., invalid JSON and unresolved code), and text augmentation (e.g., synonym replacement and reorder of words). Second, for specialized tasks in “vertical" domains, the overfitting of specific prompts occurs Ma et al. (2023) when the diversity of input instructions is rather limited. Despite the accuracy and rationality of the instruction-response pairs, LLMs tend to overfit certain patterns of the input instruction rather than truly comprehend the task. Therefore, the increase of dataset size instead reduces the generalization of trained LLMs with lower instruction following capabilities. Third, the forgetting Zhang and Wu (2024); Jin and Ren (2024b); Wang et al. (2024b) becomes a severe problem when more instruction datasets are introduced without setting a proper re-playing schedule of pre-training or previously visited instruction tuning datapoints Parmar et al. (2024); Jin et al. (2021); Ibrahim et al. (2024). The skill cultivation of a LLM on any new instruction task heavily relies on its preceding skills acquired during pre-training or previous fine-tuning. Consequently, the expansion of samples for high-level skills would “dilute" those for low-level skills and degrade performance.
@@
To sum up, with the dataset scaling up, fine-tuning with the selected subset instead of the entire set becomes a must-have strategy. To help determine the optimal selection ratio, we suggest the following three guidelines: 1) One may first develop a complex quality measure scheme that uses both indicators and human verification to estimate the noise percentage of each constituting dataset. Without lose of generality, random sampling can be performed to accelerate quality measurement. To combat noise, a lower ratio (i.e., smaller $S_{b}$) should be considered for data selection from the noisy $S$. 2) To combat overfitting, both the diversity of instruction datapoints within and across datasets should be emphasized. A higher keeping proportion should be established for datasets with diverse instruction styles, prompt formations, and response patterns, which helps improve the model’s instruction following capabilities. 3) For continual fine-tuning, datasets that share similar distributions with pre-training and previous fine-tuning datapoints should be kept to fight against forgetting. The optimal selection ratio and proportion for each dataset is built upon the meticulous and thorough analysis on each instruction dataset, and therefore case-by-case adjustment is needed. For future studies, one may investigate the automation of assessment and selection recipe to minimize the human intervention.
@@
In consideration of the cost of building human-labeled and human-verified instruction tuning datasets, methods that employ powerful LLMs like GPT4 for instruction synthesis Bradley et al. (2023); Li et al. (2023e); Xu et al. (2023a); Li et al. (2024a); Zhao et al. (2024a); Dong et al. (2024) have gained increasing attention. The synthetic instructions proliferate cost-effectively with fine-grained control of characteristics such as difficulty and style. Therefore, it is expected to witness a surge of instruction datapoints (e.g., tens or even hundreds millions) in the short future. In that case, datasets of such quantity pose a significant challenge to the robustness, efficiency, and precision of the selection methods. Previous studies like DSIM Xie et al. (2023) demonstrated that cheap approximation of features by bag-of-n-grams achieves similar performance but requires much less computing resources. For future research, one may draw inspiration from the data deduplication and filtering techniques in handling billions of pre-training tokens. Especially for the measurement of diversity, the computing of embedding-based pairwise similarity and clustering can be greatly reduced with simplified representations. In addition, the hierarchical philosophy Hmida et al. (2016); Talavera (1999); Cabezas et al. (2023); Ran et al. (2023) might be a promising approach to select data from coarse-grained to fine-grained structures. One may apply the devide-and-conquer strategy to recursively handle each subset of the instruction dataset, limiting the peak resource consumption under budget.
@@ 7.4 Scaling Up LLMs
On the other hand, under the context of data evaluation, it calls upon on rethinking of traditional machine learning techniques such as efficient optimization tricks and dimensionality reduction approaches. For example, in the assessment of loss-based datapoint influence Feldman and Zhang (2020), the exhaustive measurement on the marginal performance by moving-each-sample-out and model re-training can be simply approximated by iterative batch-wise sampling tricks with a greedy principle behind. For efficient assessment, PCA Xu et al. (2023b) and random projection Xia et al. (2024a); Park et al. (2023) are popular choices for obtaining low-rank representations of embeddings and gradients, which facilitates not only metric computation but also storing of datapoints.
@@
Recent studies on the effectiveness of instruction tuning in injecting task-specific or domain-specific knowledge into LLMs Shi and Lipani (2023); Goyal et al. (2023); Zhang et al. (2024b); Yıldız et al. (2024) show that the stand-alone instruction tuning might not be the most appropriate method. Compared with strategies like continual pre-training Cossu et al. (2024); Ke (2024); Cossu et al. (2024) and instruction modeling Lou et al. (2024); Cheng et al. (2024); Wang et al. (2022); Xu et al. (2024); Shi et al. (2024), instruction tuning counts the response sequences for loss computation without sufficient perception of instruction context. For specialized domains like medicine, finance, and laws, if the pre-trained LLMs are in lack of the prerequisite knowledge, the instruction tuning cannot properly activates the parameterized “memory" for alignment but only causes overfitting of the given prompt. In that case, the benefits of data selection are limited with poor generalizability.
Another noteworthy phenomenon in data assessment and selection studies is that due to limited budgets of computing resources, most of the experiments are performed on LLMs of small and moderate size (e.g., less than 7B) to validate the effectiveness of the quality measurement and the selection strategy. Small pre-trained LLMs, by their nature of small parameter size, are more sensitive to the instruction datasets during fine-tuning or continual learning Schick and Schütze (2020); Yıldız et al. (2024). They exhibit the most significant rates of both forggeting (old knowledge) and learning (new knowledge). In the light of such statement, small LLMs tend to sacrifice the task-irrelevant knowledge in return for rapid adaptation towards novel domains and tasks. The selected datasets by various quality measures can impose immediate effect on the parameters of small LLMs, but may weaken on those of huge ones. It remains unknown whether the same quality measurement and data selection pipeline can achieve similar performance gains on both small and large LLMs. For future research of data assessment and selection, extensive experiments are required to validate their efficiency on huge LLMs (e.g., 70B and 405B) AI@Meta (2024) and LLMs of mixture-of-experts (MoE) architectures (e.g., Mixtral 8x22B) Jiang et al. (2024a).

### v4

@@ 7.2 Unveiling the Definitions of Good Data
What signifies the most a good datapoint remains an open question.
Unfortunately, there exists no unified criteria on discriminating "good" instructions from "bad" ones. Essentially, the definitions on the general data "quality" differ from task to task and domain to domain Evans & Murshudov (2013); Flach (2012); Albalak et al. (2024). Although existing quality measurement methods can be categorized in terms of quality, diversity, and importance under the present study, they all exhibit more or less ad-hoc properties in methodology. First, studies on instruction tuning are often targeted at improving the performance of LLMs on downstream task. No matter whether these tasks are of general-purpose (e.g., common NLP tasks on leaderboard Myrzakhan et al. (2024); Wolf et al. (2019)) or domain-specific applications, such task-orientated data selection itself is only a "proxy" for exploring the underlying "quality" measurement. Especially for coreset sampling methods that directly employ the evaluation set or testing set for distribution matching or importance estimation, instructions that resemble the most to the testing set or bring about performance gains are judged as "good" data. However, such "good" data cannot be easily transferred to another LLM of completely different architecture and parameters. Each time the entire pipeline has to be enforced for a novel task, making it difficult to accumulate universally-acknowledged high-quality data for archiving. Second, each method has an individual quality evaluation system and very few of them ever tried to justify their design and interpret the philosophy behind. It is difficult to validate whether certain component of the selection pipeline can be replaced or removed for better serving a new task-of-interest.
Accordingly, further academic explorations include: 1) to present a more unified, generally applicable definitions on "good" datapoints in terms of fine-grained aspects, and 2) to improve interpretability and explanability of the selection pipeline beyond empirical design.
@@
The three aspects we used to categorize data assessment methods are actually overlapping with each other, where the "boundary" between two measuring dimension is often hard to explicitly defined. Under such circumstance, the definition of good data can be perceived as the weighted, biased mixture of quality, diversity, and importance. Existing methods are not flexible in dynamically adjusting the mixing weights to adapt to different downstream tasks. Instead, their priority order of the three dimensions is implicitly encoded into the selection of instructions. For instance,  Liu et al. (2023b) emphasizes quality and importance equally by first establishing the relative ranking of all samples in both quality and complexity. The subset is formed by consecutively selecting the top-ranked samples in sequence, with diversity intervened via ruling out heavily homogenized examples. Such hard-coded, greedy treatment to quality, diversity, and importance is not applicable to scenarios where the behavior of LLMs is expected to cater to varied preference.
@@ 7.3 Scaling Up Datasets
In the analysis of the disadvantages in exploiting the entire instruction dataset for alignment, putting aside the issue of long training time, we notice that the performance of fine-tuning the entire dataset might not be the optimal. There often exists a critical point of the best selection proportion, and such proportion varies from dataset to dataset. When more instruction datasets from diverse domains and tasks are incorporated, it becomes more difficult to nail down the best selection proportion for three main reasons. First, a large proportion of noise exists in the open datasets and few noisy samples can already cause tremendous negative impact on performance Song et al. (2022). During the pre-processing of instruction dataset, noise can be unintentionally introduced in instruction preparation (e.g., missing context or system prompt), response generation (e.g., unverified or mismatched answers), format wrapping (e.g., invalid JSON and unresolved code), and text augmentation (e.g., synonym replacement and reorder of words). Second, for specialized tasks in "vertical" domains, the overfitting of specific prompts occurs Ma et al. (2023) when the diversity of input instructions is rather limited. Despite the accuracy and rationality of the instruction-response pairs, LLMs tend to overfit certain patterns of the input instruction rather than truly comprehend the task. Therefore, the increase of dataset size instead reduces the generalization of trained LLMs with lower instruction following capabilities. Third, the forgetting Zhang & Wu (2024); Jin & Ren (2024b); Wang et al. (2024b) becomes a severe problem when more instruction datasets are introduced without setting a proper re-playing schedule of pre-training or previously visited instruction tuning datapoints Parmar et al. (2024); Jin et al. (2021); Ibrahim et al. (2024). The skill cultivation of a LLM on any new instruction task heavily relies on its preceding skills acquired during pre-training or previous fine-tuning. Consequently, the expansion of samples for high-level skills would "dilute" those for low-level skills and degrade performance.
@@
To sum up, with the dataset scaling up, fine-tuning with the selected subset instead of the entire set becomes a must-have strategy. To help determine the optimal selection ratio, we suggest the following three guidelines: 1) One may first develop a complex quality measure scheme that uses both indicators and human verification to estimate the noise percentage of each constituting dataset. Without lose of generality, random sampling can be performed to accelerate quality measurement. To combat noise, a lower ratio (i.e., smaller $S_{b}$) should be considered for data selection from the noisy $S$. 2) To combat overfitting, both the diversity of datapoints within and across datasets should be emphasized. A higher keeping proportion should be established for datasets with diverse instruction styles, prompt formations, and response patterns, which helps improve the model’s instruction following capabilities. 3) For continual fine-tuning, datasets that share similar distributions with pre-training and previous fine-tuning datapoints should be kept to fight against forgetting. The optimal selection ratio and proportion for each dataset is built upon the meticulous and thorough analysis on each instruction dataset, and therefore case-by-case adjustment is needed. For future studies, one may investigate the automation of assessment and selection recipe to minimize the human intervention.
@@
In consideration of the cost of building human-labeled and human-verified instruction tuning datasets, methods that employ powerful LLMs like GPT4 for instruction synthesis Bradley et al. (2023); Li et al. (2023f); Xu et al. (2023a); Li et al. (2024a); Zhao et al. (2024a); Dong et al. (2024) have gained increasing attention. The synthetic instructions proliferate cost-effectively with fine-grained control of characteristics such as difficulty and style. Therefore, it is expected to witness a surge of datapoints (e.g., tens or even hundreds millions) in the short future. In that case, datasets of such quantity pose a significant challenge to the robustness, efficiency, and precision of the selection methods. Previous studies like DSIM Xie et al. (2023) demonstrated that cheap approximation of features by bag-of-n-grams achieves similar performance but requires much less computing resources. For future research, one may draw inspiration from the data deduplication and filtering techniques in handling billions of pre-training tokens. Especially for the measurement of diversity, the computing of embedding-based pairwise similarity and clustering can be greatly reduced with simplified representations. In addition, the hierarchical philosophy Hmida et al. (2016); Talavera (1999); Cabezas et al. (2023); Ran et al. (2023) might be a promising approach to select data from coarse-grained to fine-grained structures. One may apply the devide-and-conquer strategy to recursively handle each subset of the instruction dataset, limiting the peak resource consumption under budget.
@@ 7.4 Scaling Up LLMs
On the other hand, under the context of data evaluation, it calls upon on rethinking of traditional machine learning techniques such as efficient optimization tricks and dimensionality reduction approaches. For example, in the assessment of loss-based datapoint influence Feldman & Zhang (2020), the exhaustive measurement on the marginal performance by moving-each-sample-out and model re-training can be simply approximated by iterative batch-wise sampling tricks with a greedy principle behind. For efficient assessment, PCA Xu et al. (2023b) and random projection Xia et al. (2024a); Park et al. (2023) are popular choices for obtaining low-rank representations of embeddings and gradients, which facilitates not only metric computation but also storing of datapoints.
@@
Recent studies on the effectiveness of instruction tuning in injecting task-specific or domain-specific knowledge into LLMs Shi & Lipani (2023); Goyal et al. (2023); Zhang et al. (2024b); Yıldız et al. (2024) show that the stand-alone instruction tuning might not be the most appropriate method. Compared with strategies like continual pre-training Cossu et al. (2024); Ke (2024); Cossu et al. (2024) and instruction modeling Lou et al. (2024); Cheng et al. (2024); Wang et al. (2022); Xu et al. (2024); Shi et al. (2024), instruction tuning counts the response sequences for loss computation without sufficient perception of instruction context. For specialized domains like medicine, finance, and laws, if the pre-trained LLMs are in lack of the prerequisite knowledge, the instruction tuning cannot properly activates the parameterized "memory" for alignment but only causes overfitting of the given prompt. In that case, the benefits of data selection are limited with poor generalizability.
Another noteworthy phenomenon in data assessment and selection studies is that due to limited budgets of computing resources, most of the experiments are performed on LLMs of small and moderate size (e.g., less than 7B) to validate the effectiveness of the quality measurement and the selection strategy. Small pre-trained LLMs, by their nature of small parameter size, are more sensitive to the instruction datasets during fine-tuning or continual learning Schick & Schütze (2020); Yıldız et al. (2024). They exhibit the most significant rates of both forggeting (old knowledge) and learning (new knowledge). In the light of such statement, small LLMs tend to sacrifice the task-irrelevant knowledge in return for rapid adaptation towards novel domains and tasks. The selected datasets by various quality measures can impose immediate effect on the parameters of small LLMs, but may weaken on those of huge ones. It remains unknown whether the same quality measurement and data selection pipeline can achieve similar performance gains on both small and large LLMs. For future research of data assessment and selection, extensive experiments are required to validate their efficiency on huge LLMs (e.g., 70B and 405B) AI@Meta (2024) and LLMs of mixture-of-experts (MoE) architectures (e.g., Mixtral 8x22B) Jiang et al. (2024a).
@@
7.5 Validating the Bias and Fairness of Instruction-tuned LLMs
As mentioned in the Sec. 1.3, most existing data selection methods fail to specifically study the dataset bias for consideration of fairness. It is noted that for domain-specific tasks such as solving programming and maths problems, the data bias might not be that severe. However, for general question-answering, recommendation, and creative writing tasks, even ChatGPT could produce biased answers that are intolerable in domains like education Doan et al. (2024); Zhang et al. (2023b); Chisca et al. (2024); Gao et al. (2024). To improve the fairness of LLMs, it is necessary to take the measurement of data bias into serious consideration. Future work includes: 1) the integration of embedding-based (e.g., WEAT Caliskan et al. (2017), SEAT May et al. (2019)), probability-based (e.g., DisCo Webster et al. (2020), CrowS-Pairs Score Nangia et al. (2020)), and generated texts-based (e.g., Co-Occurrence Bias Score Bordia & Bowman (2019), Full Gen Bias Smith et al. (2022)) techniques for fairness evaluation of datapoints Gallegos et al. (2024), and 2) the construction of bias and fairness benchmarks Jiao et al. (2024) under various domains and tasks as supplementary kits to the current evaluation pipeline that only focuses on testing the capability of LLMs.